# A parallel framework for graphical optimal transport

### v1

@@ 4.1 Algorithm
            Round $B_{(j,k)}$ by Algorithm 2 in Altschuler et al. (2017) such that $\widehat{B}_{(j,k)}\in\Pi(\bar{q}_{j}(B),B^{\top}_{(j,k)}{\bf 1})$ for $\forall k\in N(j)$, where $\bar{q}_{j}(B)=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}\mathbf{1}$.
@@
Proposition 1.
The updates in Theorem 1 can be equivalently expressed as the updates 36-37 in Algorithm 3 together with the variables 35.
The updates of the dual variables, as specified in 36-37, are independent within the same group $S$, facilitating simultaneous updates in parallel. Provided there is sufficient memory to store all variables, concurrent updating is feasible.
In each iteration Algorithm 3 updates the variables corresponding to one of the partitions $S_{1}$ or $S_{2}$. This update guarantees that the constraints are matched exactly for the nodes in this partition. However, this means that after any finite number of iterations of Algorithm 3 the marginal constraints will not be matched for the nodes in the partition that has not been updated in the last iteration. To address this, we adopt a strategy akin to the ones in Altschuler et al. (2017), Fan et al. (2022a) and employ a Rounding Algorithm 4. This ensures the generation of transport plans that adhere to all stipulated constraints in 30. The complexity of the rounding process is ${\mathcal{O}}(|E|d^{2})$. Algorithm 4 differs from the method in Fan et al. (2022a) in that it requires modification of the transport plans across all edges.
@@ 4.2 Complexity analysis
In this section, we provide a rigorous demonstration that our Algorithm 3, when used in conjunction with the Rounding Algorithm 4, requires $T=\widetilde{\mathcal{O}}\left({|E|^{2}}/{\delta^{2}}\right)$ iterations to achieve a $\delta$-approximation of the problem 28. First, we will illustrate that, through the specific mass normalization technique employed in 37, the transport plan maintains unit mass in each iteration.
Lemma 1.
@@
We now define the variable $\lambda^{t}_{(j,k)}=\epsilon\log(u^{t}_{(j,k)})$ for $(j,k)\in E$, and $u^{t}_{(j,k)}$ are the iterates in Algorithm 3. Lemma 2 provides bounds on the scaling variables.
Lemma 2.
@@
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq 2C_{\infty},\qquad\max_{i=1,\dots,d}(\lambda_{(j,k)}^{*})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{*})_{i}\leq 2C_{\infty}.$ (38)
The following lemma links the dual objective error in Theorem 1 to the stopping criteria in Algorithm 3.
Lemma 3.
@@
$f(u^{t},\rho^{t})-f(u^{*},\rho^{*})\leq\frac{2C_{\infty}}{\epsilon}\left(\sum_{j\in{\Gamma}}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\in{V\backslash\Gamma}}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}\right)$ (39)
@@
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})\geq\frac{1}{22|E|}\left(\sum_{j\in{\Gamma}}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\in{V\backslash\Gamma}}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}\right)^{2}.$ (40)
Unlike Kroshnin et al. (2019), our Lemma 3 and Lemma 4 apply universally for any $t$, obviating the need to consider odd and even $t$ distinctly.
Theorem 2.
@@
Let $\{B_{(j,k)}\}$ be a sequence of transportation plans on a bipartite tree. Algorithm 4 returns $\{\widehat{B}_{(j,k)}\}$ satisfying $\widehat{B}_{(j,k)}^{\top}{\bf 1}=q_{j}$ for $j\in V$, where $q_{j}$ is the marginal distribution of node $j$. Moreover, it holds that $\bigg{|}\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}\bigg{|}\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(B)-B_{(j,k)}{\bf 1}\|_{1}\right),$ where $\bar{q}_{j}(B)=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}\mathbf{1}$.
Lemma 6.
@@
Let $\{\widetilde{B}_{(j,k)}\}$ be the output of Algorithm 3 and let $\{\widehat{B}_{(j,k)}\}$ be the output of Algorithm 4 with input $\{\widetilde{B}_{(j,k)}\}$, and denote $\{B^{*}_{(j,k)}\}$ the optimal solution to the unregularized MOT problem 28. Then it holds that
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}$ (41a)
$\leq$ $2\epsilon|E|\log(d)+4C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(\widetilde{B})-\widetilde{B}_{(j,k)}{\bf 1}\|_{1}\right).$ (41b)
@@
Given a predetermined final error threshold $\delta$, we select parameters $\epsilon=\frac{\delta}{4|E|\log(d)}$ and $\delta^{\prime}=\frac{\delta}{8C_{\infty}}$ for use in Algorithm 3. This results in an output of $\widetilde{B}_{(j,k)}$, for each edge pair $(j,k)\in E$. Next, we employ Algorithm 4 to round $\widetilde{B}_{(j,k)}$ for each $(j,k)\in E$, thereby producing $\widehat{B}_{(j,k)}$. With these procedures in place, we are equipped to compute the final complexity for the coupled bi-marginal OT problem as described in 28.
Theorem 3.
@@
Algorithm 3, in conjunction with Algorithm 4, yields a $\delta$-approximate solution to the tree-structured MOT problem 28 in $T$ arithmetic operations, where $T=\mathcal{O}\left(\frac{|E|^{3}C_{\infty}^{2}d^{2}\log(d)}{\delta^{2}}\right).$
Remark 2 (Comparison of complexity with global regularization (Fan et al. 2022a)).
@@
We compare our complexity results with the ones in Fan et al. (2022a), for graph-structured MOT with global entropic regularization in terms of the total complexity and the number of iterations. i) Total Complexity: According to Theorem 2 in Fan et al. (2022a), global entropy regularization requires $\widetilde{\mathcal{O}}(d(G)|E||\Gamma|^{2}C_{\infty}^{2}d^{2}/\delta^{2})$ arithmetic operations with high probability, where $d(G)$ denotes the maximum node distance in graph $G$. Under the assumptions $|E|={\mathcal{O}}(|\Gamma|)$ and $d(G)={\mathcal{O}}(1)$, akin to the Wasserstein barycenter problem, both the global and local regularization methods offer similar complexity bounds. However, a notable distinction is that Theorem 3 offers a deterministic bound. ii) Iterations: Algorithm 3 with local regularization, as per Theorem 2 and Theorem 3, necessitates $\mathcal{O}\left(\frac{|E|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right)$ iterations, less than the requirement for global regularization at ${\mathcal{O}}\left(\frac{|E||\Gamma|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right)$. The reduced iteration count is a benefit of algorithm parallelization, albeit with a higher computational load per iteration.
5 Extension of the algorithm with local regularization to general
graphs
@@ 5 Extension of the algorithm with local regularization to general
Building on the concepts from Sec. 4, we extend our approach to OT problems with a general factor graph structure as indicated in 10. We employ a strategy that breaks down the MOT problem into interconnected smaller MOT problems, introducing local regularization for each sub-problem. To facilitate this, we define a specialized type of junction tree, elaborated in Definition 1.
Definition 4.
@@
Note that each separator set in Definition 1 is connected to exactly two cliques. On the other hand, by Definition 4, our cliques in ${\mathcal{Q}}_{S}$ can be connected to an arbitrary number of cliques, and instead, the cost-cliques in ${\mathcal{Q}}_{C}$ can only be connected to two cliques. This latter assumption is necessary for efficient parallelization of the algorithm we introduce and can be relaxed as we will explain in remark 3.
Example 3.
The junction trees in Figures 6, 7, and 8 are modified junction trees for the respective graphs. In contrast to junction trees in Figures 2, 3, and 4, each clique (shown as circles) can only be connected to two separators (shown as squares), while separator can connect to multiple cliques. Moreover, note that constraints in the respective MOT problems are indicated by gray boxes in Figures 6, 7, and 8. This is explained in more detail in the following section.
@@
[…]
@@ 5.1 Problem formulation
We structure a coupled MOT problem 42, which is equivalent to MOT problem 11, on the junction tree in Definition 4 by associating every cost term with a clique in ${\mathcal{Q}}_{C}$. Moreover, we assume that each constraint is embedded in a separator in ${\mathcal{Q}}_{S}$ as described in the following. A constraint of the optimal transport problem can act on projections of the solution tensors ${\mathbf{B}}_{c}$ onto a set of marginals. Recall that $\Gamma\subset\{1,2,\dots,J\}$ denote the set of index-sets corresponding to the constraints (see 2). Then we assume that there exists a partition of $\Gamma$ such that for each ${\boldsymbol{\gamma}}$ in the partition there is a $c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{S}$ such that ${\boldsymbol{\gamma}}\subseteq c_{\boldsymbol{\gamma}}$, and any $c_{\boldsymbol{\gamma}}$ can contain only one constrained set ${\boldsymbol{\gamma}}$. Moreover, we denote the separators with constraints as ${\mathcal{Q}}_{\Gamma}=\{c_{\boldsymbol{\gamma}}\}{\subseteq{\mathcal{Q}}_{S}}$. Note that ${\boldsymbol{\gamma}}$ may not only contain one node, for example, if $c_{\boldsymbol{\gamma}}=\{x_{1},x_{2},x_{3}\}$, we allow ${\boldsymbol{\gamma}}=\{x_{1},x_{3}\}$. Thus, we define the coupled MOT problem
$\min_{{\mathbf{B}}_{c},c\in{\mathcal{Q}}}$ $\sum_{c\in{\mathcal{Q}}_{C}}{\langle}{\bf C}_{c},{\mathbf{B}}_{c}{\rangle}$ (42)
subject to $P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{i}})=P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{j}}),\quad\text{ for each }c_{i}\in{\mathcal{Q}}_{S},\text{ and all }c_{j}\in N(c_{i}),$
@@
$P_{\boldsymbol{\gamma}}({\mathbf{B}}_{c_{\boldsymbol{\gamma}}})=\boldsymbol{\mu}_{\boldsymbol{\gamma}},\quad\text{for }c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}{\subseteq{\mathcal{Q}}_{S}},\qquad{\langle}{\mathbf{B}}_{c},\mathbf{1}{\rangle}=1,\quad\text{for }c{\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.}$
@@
The separators in ${\mathcal{Q}}_{S}$ enforce that the cost cliques are consistent with each other, and thereby play a similar role as the separator sets in Definition 1. To develop an algorithm for problem 42 we add a local regularization term for each clique that can contain a cost, i.e., for each $c\in{\mathcal{Q}}_{C}$. Therefore, we generalize the expression for the reference joint distribution in 29. For $c\in{\mathcal{Q}}_{C}$, let $c_{j},c_{k}\in{\mathcal{Q}}_{S}$ be the two neighbours, and define ${\bf M}_{c}=\bm{}{c_{j}}\otimes\bm{}{c_{k}}$, where $\bm{}{c_{i}}=\boldsymbol{\mu}_{\boldsymbol{\gamma}}\otimes\mathbf{1}$ if ${\boldsymbol{\gamma}}\subset c_{i}$ for some ${\boldsymbol{\gamma}}\subset\Gamma$, and otherwise $\bm{}{c_{i}}=\mathbf{1}$. We use the notation $\otimes\mathbf{1}$ to broadcast a tensor to a bigger tensor of appropriate size. In the following, the broadcasting dimensions will always be obvious from the context. For example, if ${c_{i}}=\{x_{1},x_{2},x_{3}\},{\boldsymbol{\gamma}}=\{x_{2}\},$ then $\bm{}{c_{i}}(x_{1},x_{2},x_{3})=\boldsymbol{\mu}_{\boldsymbol{\gamma}}(x_{2})$.
@@
Moreover, we add a mass constraint similar to 30. The regularized problem thus reads
$\min_{{\mathbf{B}}_{c},c\in{\mathcal{Q}}}$ $\sum_{c\in{\mathcal{Q}}_{C}}{\langle}{\bf C}_{c},{\mathbf{B}}_{c}{\rangle}+\epsilon{\mathcal{H}}({\mathbf{B}}_{c}|{\bf M}_{c})$ (43)
subject to $P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{i}})=P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{j}}),\quad\text{ for each }c_{i}\in{\mathcal{Q}}_{S},\text{ and all }c_{j}\in N(c_{i})$
@@
$P_{\boldsymbol{\gamma}}({\mathbf{B}}_{c_{\boldsymbol{\gamma}}})=\boldsymbol{\mu}_{\boldsymbol{\gamma}},\quad\text{for }c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}\subseteq{\mathcal{Q}}_{S},\qquad{\langle}{\mathbf{B}}_{c},\mathbf{1}{\rangle}=1,\quad\text{for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.$

### v2

@@ 3.1 Algorithm
     Round $B_{(j,k)}$ by [altschuler2017near, Algorithm 2] such that $\widehat{B}_{(j,k)}\in\Pi(\bar{q}_{j}(B),B^{\top}_{(j,k)}{\bf 1})$ for $\forall k\in N(j)$, where $\bar{q}_{j}(B)=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}\mathbf{1}$.
@@
Proposition 3.3.
The updates in Theorem 3.1 can be equivalently expressed as the updates Eq. 27-Eq. 28 in Algorithm 2 together with the variables Eq. 26.
The updates of the dual variables, as specified in Eq. 27-Eq. 28, are independent within the same group $S$, facilitating simultaneous updates in parallel. Provided there is sufficient memory to store all variables, concurrent updating is feasible.
In each iteration Algorithm 2 updates the variables corresponding to one of the partitions $S_{1}$ or $S_{2}$. This update guarantees that the constraints are matched exactly for the nodes in this partition. However, this means that after any finite number of iterations of Algorithm 2 the marginal constraints will not be matched for the nodes in the partition that has not been updated in the last iteration. To address this, we adopt a strategy akin to the ones in [altschuler2017near, fan2021complexity] and employ a Rounding Algorithm 3. This ensures the generation of transport plans that adhere to all stipulated constraints in Eq. 21. The complexity of the rounding process is ${\mathcal{O}}(|E|d^{2})$. Algorithm 3 differs from the method in [fan2021complexity] in that it requires modification of the transport plans across all edges.
@@ 3.2 Complexity analysis
In this section, we provide a rigorous demonstration that our Algorithm 2, when used in conjunction with the Rounding Algorithm 3, requires $T=\widetilde{\mathcal{O}}\left({|E|^{2}}/{\delta^{2}}\right)$ iterations to achieve a $\delta$-approximation of the problem Eq. 19. First, we will illustrate that, through the specific mass normalization technique employed in Eq. 28, the transport plan maintains unit mass in each iteration.
Lemma 3.4.
@@
[…]
Lemma 3.5.
@@
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq 2C_{\infty},\qquad\max_{i=1,\dots,d}(\lambda_{(j,k)}^{*})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{*})_{i}\leq 2C_{\infty}.$
The following lemma links the dual objective error in Theorem 3.1 to the stopping criteria in Algorithm 2.
Lemma 3.6.
@@
$f(u^{t},\rho^{t})-f(u^{*},\rho^{*})\leq\frac{2C_{\infty}}{\epsilon}\left(\sum_{j\in{\Gamma}}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\sum_{j\in{V\backslash\Gamma}}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}\right)$
@@
$f(u^{t},\rho^{t})-f(u^{t+1}\!,\rho^{t+1})\geq\frac{1}{22|E|}\!\left(\sum_{j\in{\Gamma}}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}+\!\!\sum_{j\in{V\backslash\Gamma}}\sum_{k\in N(j)}\!\!\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}\!\right)^{\!\!2}\!\!.$
Unlike [kroshnin2019complexity], our Lemma 3.6 and Lemma 3.7 apply universally for any $t$, obviating the need to consider odd and even $t$ distinctly.
Theorem 3.8.
@@
Let $\{B_{(j,k)}\}$ be a sequence of transportation plans on a bipartite tree. Algorithm 3 returns $\{\widehat{B}_{(j,k)}\}$ satisfying $\widehat{B}_{(j,k)}^{\top}{\bf 1}=q_{j}$ for $j\in V$, where $q_{j}$ is the marginal distribution of node $j$. Moreover, it holds that
$\bigg|\sum_{(j,k)\in E}{\langle}C_{(j,k)},B_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}\bigg|$
$\leq 2C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-B_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(B)-B_{(j,k)}{\bf 1}\|_{1}\right),$
where $\bar{q}_{j}(B)=\frac{1}{|N(j)|}\sum_{k\in N(j)}B_{(j,k)}\mathbf{1}$.
Lemma 3.10.
@@
Let $\{\widetilde{B}_{(j,k)}\}$ be the output of Algorithm 2 and let $\{\widehat{B}_{(j,k)}\}$ be the output of Algorithm 3 with input $\{\widetilde{B}_{(j,k)}\}$, and denote $\{B^{*}_{(j,k)}\}$ the optimal solution to the unregularized MOT problem Eq. 19. Then it holds that
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}$
$\leq$ $2\epsilon|E|\log(d)+4C_{\infty}\left(\sum_{j\in\Gamma}\|\mu_{j}-\widetilde{B}_{(j,k_{j})}{\bf 1})\|_{1}+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\|\bar{q}_{j}(\widetilde{B})-\widetilde{B}_{(j,k)}{\bf 1}\|_{1}\right).$
@@
Given a predetermined final error threshold $\delta$, we select parameters $\epsilon=\frac{\delta}{4|E|\log(d)}$ and $\delta^{\prime}=\frac{\delta}{8C_{\infty}}$ for use in Algorithm 2. This results in an output of $\widetilde{B}_{(j,k)}$, for each edge pair $(j,k)\in E$. Next, we employ Algorithm 3 to round $\widetilde{B}_{(j,k)}$ for each $(j,k)\in E$, thereby producing $\widehat{B}_{(j,k)}$. With these procedures in place, we are equipped to compute the final complexity for the coupled bi-marginal OT problem as described in Eq. 19.
Theorem 3.11.
@@
Algorithm 2, in conjunction with Algorithm 3, yields a $\delta$-approximate solution to the tree-structured MOT problem Eq. 19 in $T$ arithmetic operations, where $T=\mathcal{O}\left(\frac{|E|^{3}C_{\infty}^{2}d^{2}\log(d)}{\delta^{2}}\right).$
Remark 3.12 (Comparison of complexity with global regularization [fan2021complexity]).
@@
We compare our complexity results with the ones in [fan2021complexity], for graph-structured MOT with global entropic regularization (cf. supplementary material) in terms of the total complexity and the number of iterations. i) Total Complexity: According to Theorem 2 in [fan2021complexity], global entropy regularization requires $\widetilde{\mathcal{O}}(d(G)|E||\Gamma|^{2}C_{\infty}^{2}d^{2}/\delta^{2})$ arithmetic operations with high probability, where $d(G)$ denotes the maximum node distance in graph $G$. Under the assumptions $|E|={\mathcal{O}}(|\Gamma|)$ and $d(G)={\mathcal{O}}(1)$, akin to the Wasserstein barycenter problem, both the global and local regularization methods offer similar complexity bounds. However, a notable distinction is that Theorem 3.11 offers a deterministic bound. ii) Iterations: Algorithm 2 with local regularization, as per Theorem 3.8 and Theorem 3.11, necessitates $\mathcal{O}\left(\frac{|E|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right)$ iterations, less than the requirement for global regularization at ${\mathcal{O}}\left(\frac{|E||\Gamma|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right)$. The reduced iteration count is a benefit of algorithm parallelization, albeit with a higher computational load per iteration.
4 Generalized algorithm with local regularization to general graphs
@@
Building on the concepts from Section 3, we extend our approach to OT problems with a general factor graph structure as indicated in Eq. 8. We employ a strategy that breaks down the MOT problem into interconnected smaller MOT problems, introducing local regularization for each sub-problem. To facilitate this, we define a specialized type of junction tree, elaborated in Definition 10.1.
Definition 4.1.
@@
Note that each separator set in Definition 10.1 is connected to exactly two cliques. On the other hand, by Definition 4.1, our cliques in ${\mathcal{Q}}_{S}$ can be connected to an arbitrary number of cliques, and instead, the cost-cliques in ${\mathcal{Q}}_{C}$ can only be connected to two cliques. This latter assumption is necessary for efficient parallelization of the algorithm we introduce and can be relaxed as we will explain in Remark 4.6.
Example 4.2.
The junction trees in Figures 3, 4, and 5 are modified junction trees for the respective graphs. In contrast to junction trees for the problem with global regularization, depicted in Figures 8, 9, and 10, each clique (shown as circles) can only be connected to two separators (shown as squares), while separator can connect to multiple cliques. Moreover, note that constraints in the respective MOT problems are indicated by gray boxes in Figures 3, 4, and 5. This is explained in more detail in the following section.
@@
$\begin{subarray}{c}0,\ 1\\ J+1\end{subarray}$$\begin{subarray}{c}0,\ 2\\ J+1\end{subarray}$$\begin{subarray}{c}0,\ 3\\ J+1\end{subarray}$$\begin{subarray}{c}\!0,~J-1\!\\ J+1\end{subarray}$$\begin{subarray}{c}0,\ J\\ J+1\end{subarray}$0,J+10,J+1112233J1J-1JJ
@@ 4.1 Problem formulation
We structure a coupled MOT problem Eq. 29, which is equivalent to MOT problem Eq. 9, on the junction tree in Definition 4.1 by associating every cost term with a clique in ${\mathcal{Q}}_{C}$. Moreover, we assume that each constraint is embedded in a separator in ${\mathcal{Q}}_{S}$ as described in the following. A constraint of the optimal transport problem can act on projections of the solution tensors ${\mathbf{B}}_{c}$ onto a set of marginals. Recall that $\Gamma\subset\{1,2,\dots,J\}$ denote the set of index-sets corresponding to the constraints (see Eq. 2). Then we assume that there exists a partition of $\Gamma$ such that for each $\bm{\gamma}$} in the partition there is a $c_{\bm{\gamma}}\in{\mathcal{Q}}_{S}$ such that ${\bm{\gamma}}\subseteq c_{\bm{\gamma}}$, and any $c_{\bm{\gamma}}$ can contain only one constrained set $\bm{\gamma}$}. Moreover, we denote the separators with constraints as ${\mathcal{Q}}_{\Gamma}=\{c_{\bm{\gamma}}\}{\subseteq{\mathcal{Q}}_{S}}$. Note that $\bm{\gamma}$} may not only contain one node, for example, if $c_{\bm{\gamma}}=\{x_{1},x_{2},x_{3}\}$, we allow ${\bm{\gamma}}=\{x_{1},x_{3}\}$. Thus, we define the coupled MOT problem
(29) $\min_{{\mathbf{B}}_{c},c\in{\mathcal{Q}}}$ $\sum_{c\in{\mathcal{Q}}_{C}}{\langle}{\bf C}_{c},{\mathbf{B}}_{c}{\rangle}$
s.t. $P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{i}})=P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{j}}),\quad\text{ for each }c_{i}\in{\mathcal{Q}}_{S},\text{ and all }c_{j}\in N(c_{i}),$
@@
$P_{\bm{\gamma}}({\mathbf{B}}_{c_{\bm{\gamma}}})=\bm{\mu}_{\bm{\gamma}},\quad\text{for }c_{\bm{\gamma}}\in{\mathcal{Q}}_{\Gamma}{\subseteq{\mathcal{Q}}_{S}},\qquad{\langle}{\mathbf{B}}_{c},\mathbf{1}{\rangle}=1,\quad\text{for }c{\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.}$
@@
The separators in ${\mathcal{Q}}_{S}$ enforce that the cost cliques are consistent with each other, and thereby play a similar role as the separator sets in Definition 10.1. To develop an algorithm for problem Eq. 29 we add a local regularization term for each clique that can contain a cost, i.e., for each $c\in{\mathcal{Q}}_{C}$. Therefore, we generalize the expression for the reference joint distribution in Eq. 20. For $c\in{\mathcal{Q}}_{C}$, let $c_{j},c_{k}\in{\mathcal{Q}}_{S}$ be the two neighbours, and define ${\bf M}_{c}=\bm{}{c_{j}}\otimes\bm{}{c_{k}}$, where $\bm{}{c_{i}}=\bm{\mu}_{\bm{\gamma}}\otimes\mathbf{1}$ if ${\bm{\gamma}}\subset c_{i}$ for some ${\bm{\gamma}}\subset\Gamma$, and otherwise $\bm{}{c_{i}}=\mathbf{1}$. We use the notation $\otimes\mathbf{1}$ to broadcast a tensor to a bigger tensor of appropriate size. In the following, the broadcasting dimensions will always be obvious from the context. For example, if ${c_{i}}=\{x_{1},x_{2},x_{3}\},{\bm{\gamma}}=\{x_{2}\},$ then $\bm{}{c_{i}}(x_{1},x_{2},x_{3})=\bm{\mu}_{\bm{\gamma}}(x_{2})$.
@@
Moreover, we add a mass constraint similar to Eq. 21. The regularized problem thus reads
(30) $\min_{{\mathbf{B}}_{c},c\in{\mathcal{Q}}}$ $\sum_{c\in{\mathcal{Q}}_{C}}{\langle}{\bf C}_{c},{\mathbf{B}}_{c}{\rangle}+\epsilon{\mathcal{H}}({\mathbf{B}}_{c}|{\bf M}_{c})$
s.t. $P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{i}})=P_{{c_{i}}\cap{c_{j}}}({\mathbf{B}}_{c_{j}}),\quad\text{ for each }c_{i}\in{\mathcal{Q}}_{S},\text{ and all }c_{j}\in N(c_{i})$
@@
$P_{\bm{\gamma}}({\mathbf{B}}_{c_{\bm{\gamma}}})=\bm{\mu}_{\bm{\gamma}},\quad\text{for }c_{\bm{\gamma}}\in{\mathcal{Q}}_{\Gamma}\subseteq{\mathcal{Q}}_{S},\qquad{\langle}{\mathbf{B}}_{c},\mathbf{1}{\rangle}=1,\quad\text{for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}.$